An angle is 275 degrees less than four times its complement, what is the value of the angles?
Answer by JulietG(1812)

Angles complement each other at 90 degrees.
Let A be the first angle, and B be the second.
A + B = 90
A = 4B - 275
Substitute the value of A from the second equation into the first.
A + B = 90
(4B-275) + B = 90
Add 275 to each side
5B = 365
Divide each side by 5
B = 73 degrees
If Angle B is 73 degrees, then Angle A is 17 degrees (90-73).

Prove it:
An angle is 275 degrees less than four times its complement
An angle (17 degrees) is 275 less than 4*73
4*73=292
292-275=17